OpenStudy (ledah):

Simplifying radicals: 8 times the square root of 63 x to the fifth power

OpenStudy (pfenn1):

Is this the expression? \[8\sqrt{63x^5}\]

OpenStudy (campbell_st):

so its \[8\sqrt{63x^5}\] break it down into smaller problems \[8\sqrt{9 \times 7 \times x^4\times x} = 8\times \sqrt{9}\times \sqrt{x^4}\times \sqrt{7x}\] Hopefully this helpsyou to get the answer

OpenStudy (ledah):

@pfenn1 yeah thats the expression

OpenStudy (pfenn1):

If that expression is correct, just follow what @campbell_st indicated. It should help you see what to do.